# The finite groups with three automorphism orbits

### v1

@@ 3. Exterior Square of Transitive Linear Groups
Note that $G$ is isomorphic one of groups in Theorem 3.1. Using Magma [5], we obtain that if $G$ is one of the groups listed in Theorem 3.1 (iii) and (iv), then $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W=1$, as in case (i). Thus, we focus on groups $G$ with $G_{0}:=G^{(\infty)}$ isomorphic to one of ${\mathrm{SL}}(m,p^{f})$ ($m\geqslant 3$ and $n=mf$) and ${\mathrm{Sp}}(2m,p^{f})$ ($(m,p^{f})\neq(1,3)$ and $n=2mf$).
Assume that $G_{0}\cong{\mathrm{SL}}(m,p^{f})$ with $m\geqslant 3$ and $n=mf$. Then case (ii) holds by Lemma 3.5 (ii). Assume that $G_{0}\cong{\mathrm{Sp}}(2m,p^{f})$ with $(m,p^{f})\neq(1,3)$ and $n=2mf$. Lemma 3.7 (iv) deduces case (iii), and we complete the proof. ∎
@@ 4. $3$-orbit $p$-groups for odd prime $p$
Let $N$ be a finite non-abelian $3$-orbit $p$-group with $p$ an odd prime. Assume that $N/\Phi(N)\cong{\mathbb{Z}}_{p}^{n}$. Then $N$ is a quotient group ${\mathcal{H}}/W$ where ${\mathcal{H}}:={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Identify $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$ with ${\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$ and ${\mathbf{Z}}({\mathcal{H}})/W$, respectively. Let $V:=N/{\mathbf{Z}}(N)$ and $M:={\mathbf{Z}}(N)$. Recall Lemmas 2.8 (iii) and 2.10 that
@@
    $G:={\mathrm{Aut}}(N)^{V}$ is the maximal subgroup of $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)$ in the sense that $W^{G}=W$; and
@@ 4.1. ${\mathrm{SL}}(3,q)\lhd G$
Assume that $G_{0}=G^{(\infty)}\cong{\mathrm{SL}}(3,q)$ with $q=p^{f}$. We will show that such $3$-orbit group is uniquely determined by its order, and it is isomorphic to $A_{p}(n,\theta)$ with $|\theta|=3$ given in Example 2.4, as in the 6-th row of Table B.
@@
    If $N$ is a $3$-orbit $p$-group such that $|N|=|{\mathbf{Z}}(N)|^{2}$ then $|{\mathbf{Z}}(N)|=q^{3}$ for some $q=p^{f}$ and ${\mathrm{SL}}(3,q)\lhd{\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$.
@@
    Let $N$ be the $3$-orbit $p$-group with $|N|=|{\mathbf{Z}}(N)|^{2}=q^{6}$ and $q=p^{f}$, then ${\mathrm{Aut}}(N)\cong p^{9f^{2}}{:}{\Gamma\mathrm{L}}(3,q)$.
@@
Assume that $N$ is a $3$-orbit $p$-group such that $|N|=|{\mathbf{Z}}(N)|^{2}$. By Theorem 2.16, ${\mathrm{Aut}}(N)$ must be non-solvable. Let $V:=N/{\mathbf{Z}}(N)$, $M:={\mathbf{Z}}(N)$ and $G:={\mathrm{Aut}}(N)^{V}$. Then $G$ acts transitively on non-zero vectors of both $V$ and $M$ by Lemma 2.8 (iii); and $M$ is isomorphic to $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ as ${\mathbb{F}}_{p}G$-modules by Lemma 2.8 (iv). Since $|M|=|V|$, we obtain that $G^{(\infty)}\cong{\mathrm{SL}}(3,q)$ and $|M|=|V|=q^{3}$ by Theorem A, as in part (i). Let ${\mathcal{H}}:={\mathcal{H}}_{n,p}$ with $n=3f$, and let $V:={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. Then $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$ and ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ as ${\mathbb{F}}_{p}X$-modules by Lemma 2.10 (i). Let $G\leqslant X$ such that $G\cong{\Gamma\mathrm{L}}(3,q)$. Then ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ can be decomposed into $A\oplus B$ as ${\mathbb{F}}_{p}G$-modules such that $G$ acts transitively on non-zero vectors of ${\mathbf{Z}}({\mathcal{H}})/B\cong A$ by Lemma 3.5 (ii). Let $N:={\mathcal{H}}/B$. Then $G\leqslant{\mathrm{Aut}}(N)^{V}$ by Lemma 2.10 (ii), and $G$ acts transitively on non-identity elements of both $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$. Hence, $N$ is a $3$-orbit group by Lemma 2.8 (iii). This proves the existence in part (ii).
@@
Now we prove the uniqueness in part (ii) by assuming that $N_{1}={\mathcal{H}}/W_{1}$ is another desired $3$-orbit $p$-group. Then part (i) deduces that ${\mathrm{SL}}(3,q)\lhd G_{1}:={\mathrm{Aut}}(N_{1})^{V}$. Note that there exists $x\in X$ such that $G_{1}^{x}\leqslant G$, and then $N_{1}\cong{\mathcal{H}}/W^{x}$ by Lemma 2.12. Thus, we may assume that $G_{1}\leqslant G$. Let $G_{0}=G^{(\infty)}=G_{1}^{(\infty)}\cong{\mathrm{SL}}(3,q)$. Then $W_{1}$ and $B$ are ${\mathbb{F}}_{p}G_{0}$-submodules of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ such that $G_{0}$ acts transitively on non-zero vectors of the corresponding quotient modules. By Lemma 3.5, we obtain that $W_{1}=B$, and therefore $N_{1}=N$.
@@
Let $N={\mathcal{H}}/B$ as defined above. Then ${\mathrm{SL}}(3,q)\lhd{\mathrm{Aut}}(N)^{V}$ by part (i), and hence $G={\mathrm{Aut}}(N)^{V}\leqslant{\Gamma\mathrm{L}}(3,q)$. Note that $B^{g}=B$ for any $g\in{\Gamma\mathrm{L}}(3,q)$ by Lemma 3.5 (ii). Thus, we obtain that $G={\Gamma\mathrm{L}}(3,q)$ and ${\mathrm{Aut}}(N)\cong p^{n^{2}}{:}{\Gamma\mathrm{L}}(3,q)$ by Lemma 2.8 (i) and 2.11, as in part (iii). ∎
@@
Recall that $|A_{p}(n,\theta)|=|{\mathbf{Z}}(A_{p}(n,\theta))|^{2}=p^{2n}$ given in Example 2.4. Observe that ${\Gamma\mathrm{L}}(3,p^{n/3})$ contains a Singer cycle, the cyclic subgroup of order $p^{n}-1$, and ${\mathrm{Aut}}(A_{p}(n,\theta))$ also has an automorphism of order $p^{n}-1$ given in Proposition 2.5 (ii). This is the key to prove $A_{p}(n,\theta)$ is a $3$-orbit group for $|\theta|=3$. We need the following lemma before proving the isomorphism.
@@
Let ${\mathrm{F}}:={\mathbb{F}}_{p}$ and ${\mathrm{E}}:={\mathbb{F}}_{q}$. It is clear that $\Lambda^{2}_{{\mathrm{F}}}(V)$ is a direct sum of irreducible ${\mathrm{F}}G$-submodules by Maschke’s Theorem as $p\nmid|G|$. Note that $G\cong{\mathrm{GL}}(1,p^{n})$. Let $U$ be the natural ${\mathrm{E}}G$-module, and let $\phi$ be a generator of ${\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$. Then clearly that $U\cong V$ as ${\mathrm{F}}G$-modules. By Lemma 3.2, $\Lambda_{\mathrm{F}}^{2}(V)=\oplus_{i=1}^{\lfloor\frac{n}{2}\rfloor}B_{j}$ where
  1. (i)
    $B_{j}\cong U\otimes_{{\mathrm{E}}}U^{\phi^{j}}$ as ${\mathrm{F}}G$-modules for $1\leqslant j<\frac{n}{2}$;
  2. (ii)
    $2\cdot B_{n/2}\cong U\otimes_{{\mathrm{E}}}U^{\phi^{n/2}}$ as ${\mathrm{F}}G$-modules, when $n$ is even.
We claim that $B_{j}$’s are irreducible ${\mathrm{F}}G$-modules. Note that the kernel of $G$ on $B_{j}$ has order $\mathrm{gcd}(p^{j}+1,p^{n}-1)$. When $n$ is even, the induced permutation group of $G$ on $B_{n/2}$ is isomorphic to ${\mathbb{Z}}_{p^{n/2}-1}$. It follows that $B_{n/2}$ is an irreducible ${\mathbb{F}}_{p}G$-module as $\dim B_{n/2}=\frac{n}{2}$. Now we assume that $1\leqslant j<\frac{n}{2}$ and $n\geqslant 3$. When $n=6$ and $p=2$, it is not hard to see that both $B_{1}$ and $B_{2}$ are irreducible ${\mathrm{F}}G$-modules. When $(n,p)\neq(6,2)$ and $n\geqslant 3$, there exists a prime divisor $r$ of $p^{n}-1$ such that $r\mid p^{n}-1$ and $r\nmid p^{k}-1$ for any $k<n$ by Zsigmondy’s Theorem [33]. Then the order of the induced permutation group of $G$ on $B_{j}$ is divisible by $r$, and hence $B_{j}$ is an irreducible ${\mathbb{F}}_{p}G$-module as $\dim B_{j}=n$.
Since $\dim B_{n/2}\neq\dim B_{j}$ for $21\leqslant i\neq j<n/$ 2, we only need to show that $B_{i}\ncong B_{j}$ for $21\leqslant i\neq j<n/$ 2. If $B_{i}\cong B_{j}$ then $B_{i}^{\mathrm{E}}\cong B_{j}^{\mathrm{E}}$, which yields that $U\otimes_{\mathrm{E}}U^{\phi^{i}}\cong(U\otimes_{\mathrm{E}}U^{\phi^{j}})^{\phi^{k}}$ for some $k$. Let $g$ be a generator of $G$, and let $\lambda$ be the character of $g$ on $U$. Then $\lambda$ is a generator of ${\mathrm{E}}^{\times}$. Note that the character of $g$ on $U\otimes_{\mathrm{E}}U^{\phi^{i}}$ and $(U\otimes_{\mathrm{E}}U^{\phi^{i}})^{\phi^{k}}$ are $\lambda^{1+p^{i}}$ and $\lambda^{(1+p^{j})p^{k}}$, respectively. Then we have
@@
Hence, the characters of $g$ on $U\otimes_{\mathrm{E}}U^{\phi^{i}}$ and $(U\otimes_{\mathrm{E}}U^{\phi^{i}})^{\phi^{k}}$ cannot be equal. Therefore, $B_{j}$’s are pairwise non-isomorphic irreducible summands of $\Lambda^{2}_{{\mathrm{F}}}(V)$. ∎
We remark that the character of ${\mathrm{F}}G$-module $V$ is equal to the character of $V^{\mathrm{E}}$ for any extension field ${\mathrm{E}}$ of ${\mathrm{F}}$. Let $U$ be an ${\mathrm{E}}G$-module such that $U\cong V$ as ${\mathrm{F}}G$-modules. Lemma 3.2 (i) deduces that
$\chi_{V}(g)=\mathrm{Tr}_{{\mathrm{E}}/{\mathrm{F}}}\chi_{U}(g)=\sum_{i=1}^{f}\chi_{U}(g)^{\phi^{i}},$
where $\chi_{V}$ and $\chi_{U}$ are characters of $V$ and $U$, respectively, and $\langle\phi\rangle={\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$ with $[{\mathrm{E}},{\mathrm{F}}]=f$. Now we are ready to get the 6-th row of Table 1.
@@
Let $N:=A_{p}(n,\theta)$ be defined in Example 2.4 such that $3\mid n$, $|\theta|=3$ and $p$ is an odd prime. Then $N$ is a $3$-orbit group and ${\mathrm{Aut}}(N)\cong p^{n^{2}}{:}{\Gamma\mathrm{L}}(3,p^{n/3})$.
@@
Note that $N$ is a special $p$-group of exponent $p$. We assume that $N={\mathcal{H}}/W$ with ${\mathcal{H}}={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Let $V:={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})=N/{\mathbf{Z}}(N)$, and let $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$.

### v2

@@ 3. Exterior Square of Transitive Linear Groups
Note that $G$ is isomorphic one of the groups in Theorem 3.1. Using Magma [7], we obtain that if $G$ is one of the groups listed in Theorem 3.1 (iii) and (iv), then $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W=1$, as in case (i). If $G^{(\infty)}\cong{\mathrm{SL}}(m,q)$ with $m\geqslant 3$, $q=p^{f}$ and $n=mf$, then case (ii) holds by Lemma 3.2. Assume that $G_{0}\cong{\mathrm{Sp}}(2m,p^{f})$ with $(m,p^{f})\neq(1,3)$ and $n=2mf$. Lemma 3.3 implies case (iii). ∎
@@ 4. $3$-orbit $p$-groups for odd prime $p$
Let $N$ be a finite non-abelian $3$-orbit $p$-group with $p$ an odd prime. Assume that $N/\Phi(N)\cong{\mathbb{Z}}_{p}^{n}$. Then $N$ is a quotient group ${\mathcal{H}}/W$ where ${\mathcal{H}}={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Identify $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$ with ${\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$ and ${\mathbf{Z}}({\mathcal{H}})/W$, respectively. Let $V=N/{\mathbf{Z}}(N)$ and $M={\mathbf{Z}}(N)$. Recall Lemmas 2.8 (iii) and 2.10 that
@@
    $G={\mathrm{Aut}}(N)^{V}$ is the maximal subgroup of ${\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)$ in the sense that $W^{G}=W$; and
@@ 4.1. ${\mathrm{SL}}(3,q)\lhd G$
Assume that $G_{0}=G^{(\infty)}\cong{\mathrm{SL}}(3,q)$ with $q=p^{f}$. We will show that such $3$-orbit group is uniquely determined by its order, and it is isomorphic to $A_{p}(n,\theta)$ with $|\theta|=3$ given in Example 2.4, as in the 6-th row of Table 1.
@@
    If $N$ is a $3$-orbit $p$-group such that $|N|=|{\mathbf{Z}}(N)|^{2}$, then $|{\mathbf{Z}}(N)|=q^{3}$ for some $q=p^{f}$ and ${\mathrm{SL}}(3,q)\lhd{\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$.
@@
    Let $N$ be the $3$-orbit $p$-group with $|N|=|{\mathbf{Z}}(N)|^{2}=q^{6}$ and $q=p^{f}$. Then ${\mathrm{Aut}}(N)\cong p^{9f^{2}}{:}{\Gamma\mathrm{L}}(3,q)$.
@@
Assume that $N$ is a $3$-orbit $p$-group such that $|N|=|{\mathbf{Z}}(N)|^{2}$. Then ${\mathrm{Aut}}(N)$ is non-solvable by Theorem 2.16. Let $V=N/{\mathbf{Z}}(N)$, $M={\mathbf{Z}}(N)$ and $G={\mathrm{Aut}}(N)^{V}$. Lemma 2.8 yields that $G$ acts transitively on non-zero vectors of $V$ and $M$ and $M\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ for some ${\mathbb{F}}_{p}G$-submodule $W$ of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$. Hence, we have that $G^{(\infty)}\cong{\mathrm{SL}}(3,q)$ with $|M|=|V|=q^{3}$ by Theorem A, as in part (i).
@@
Let ${\mathcal{H}}={\mathcal{H}}_{n,p}$ with $n=3f$, and let $V={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. Then $X={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$ and ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ as ${\mathbb{F}}_{p}X$-modules by Lemma 2.10 (i). Let $G\leqslant X$ such that $G\cong{\Gamma\mathrm{L}}(3,q)$. Lemma 3.2 shows that there exists an ${\mathbb{F}}_{p}G$-submodule $W$ of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ such that $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ is ${\mathbb{F}}_{p}G^{(\infty)}$-isomorphic to the dual of $V$. We identify $W$ as a subgroup of ${\mathbf{Z}}({\mathcal{H}})$. Then $G\leqslant{\mathrm{Aut}}(N)^{V}$ by Lemma 2.10 (ii). It follows that $G$ acts transitively on non-identity elements of both $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$. Hence, $N$ is a $3$-orbit group by Lemma 2.8 (iii). This proves the existence in part (ii). Now we prove the uniqueness in part (ii) by assuming that $N_{1}={\mathcal{H}}/W_{1}$ is another desired $3$-orbit $p$-group. Then part (i) implies that ${\mathrm{SL}}(3,q)\lhd{\mathrm{Aut}}(N_{1})^{V}$. Denote ${\mathrm{Aut}}(N_{1})$ by $G_{1}$. Note that there exists $x\in X$ such that $(G_{1})^{(\infty)}$ is conjugate to $G^{(\infty)}$ by $x$, and hence $G_{1}^{x}\leqslant G$. Note that $W_{1}^{x}$ and $W$ are ${\mathbb{F}}_{p}G_{1}^{x}$-submodules such that $G_{1}^{x}$ acts transitively on the non-zero vectors of the corresponding factor modules. Lemma 3.2 yields that $W_{1}^{x}=W$. By 2.12, we have that
$N_{1}={\mathcal{H}}/W_{1}\cong{\mathcal{H}}/W_{1}^{x}={\mathcal{H}}/W\cong N.$
@@
Let $N={\mathcal{H}}/W$ as defined above. Then ${\Gamma\mathrm{L}}(3,q)\cong G\leqslant{\mathrm{Aut}}(N)^{V}$ as given in the proof of the existence in part (ii). By part (i), we have that ${\mathrm{SL}}(3,q)\lhd{\mathrm{Aut}}(N)^{V}$. It follows that $G={\mathrm{Aut}}(N)^{V}$, and hence ${\mathrm{Aut}}(N)\cong p^{n^{2}}{:}{\Gamma\mathrm{L}}(3,q)$ by Lemmas 2.8 (i) and 2.11, as in part (iii). ∎
@@
Recall that $|A_{p}(n,\theta)|=|{\mathbf{Z}}(A_{p}(n,\theta))|^{2}=p^{2n}$ given in Example 2.4. We aim to prove $A_{p}(n,\theta)$ is a $3$-orbit group when $3\mid n$ and $|\theta|=3$. Observe that ${\Gamma\mathrm{L}}(3,p^{n/3})$ contains a Singer cycle, the cyclic subgroup of order $p^{n}-1$, and ${\mathrm{Aut}}(A_{p}(n,\theta))$ also has an automorphism of order $p^{n}-1$ given in Proposition 2.5 (ii). We need the following lemma before proving the isomorphism.
@@
Let ${\mathrm{E}}={\mathbb{F}}_{q}$ with $q=p^{n}-1$, and let $U$ be the natural ${\mathbb{F}}_{q}G$-module. Clear that ${\mathrm{End}}_{{\mathbb{F}}_{p}G}(V)\cong{\mathrm{E}}$. We may assume that $U\cong V$ as ${\mathbb{F}}_{p}G$-modules. Notice that $\Lambda_{{\mathbb{F}}_{q}G}^{2}(U)=1$. Let $\phi$ be the Frobenius automorphism $x\mapsto x^{p}$ in ${\mathrm{E}}$. With the decomposition given in (1), we have that
$\Lambda^{2}_{{\mathbb{F}}_{p}}(V)\cong\bigoplus_{i=1}^{\lfloor\frac{n}{2}\rfloor}B_{j},$
where, as ${\mathbb{F}}_{p}G$-modules, $B_{j}\cong U\otimes_{{\mathrm{E}}}U^{\phi^{j}}$ for $1\leqslant j<\frac{n}{2}$; and $2\cdot B_{\frac{n}{2}}\cong U\otimes_{{\mathrm{E}}}U^{\phi^{n/2}}$. It suffices to show that each $B_{j}$ is an irreducible ${\mathbb{F}}_{p}G$-module and $B_{i}\ncong B_{j}$ if $i\neq j$.
First, we show that each $B_{j}$ is irreducible. Note that the kernel of $G$ on $B_{j}$ has order $\gcd(p^{j}+1,p^{n}-1)$. Then, when $n$ is even, the permutation image of $G$ on $B_{\frac{n}{2}}$ is isomorphic to ${\mathbb{Z}}_{p^{\frac{n}{2}}-1}$, which is irreducible on $B_{\frac{n}{2}}$. Assume that $j<\frac{n}{2}$ with $n\geqslant 3$. If $(n,p)=(6,2)$, it is not hard to check that $B_{1}$ and $B_{2}$ are non-isomorphic irreducible ${\mathrm{F}}G$-modules. If $(n,p)\neq(6,2)$ and $n\geqslant 3$, there exists a prime divisor $r$ of $p^{n}-1$ such that $r\mid p^{n}-1$ and $r\nmid p^{k}-1$ for any $k<n$ by Zsigmondy’s Theorem [39]. Then the order of the induced permutation group of $G$ on $B_{j}$ is divisible by $r$, and hence $B_{j}$ is an irreducible ${\mathbb{F}}_{p}G$-module as $\dim B_{j}=n$.
It is clear that $B_{\frac{n}{2}}\ncong B_{j}$ for $1\leqslant j<\frac{n}{2}$ as they have different dimensions. We only need to show that $B_{i}\ncong B_{j}$ for $21\leqslant i\neq j<n/$ 2. Note that $B_{i}\cong B_{j}$ if and only if $U\otimes_{\mathrm{E}}U^{\phi^{i}}\cong(U\otimes_{\mathrm{E}}U^{\phi^{j}})^{\phi^{k}}$ for some $k$. Let $g$ be a generator of $G$, and let $\lambda$ be the character of $g$ on $U$. Suppose that $U\otimes_{\mathrm{E}}U^{\phi^{i}}\cong(U\otimes_{\mathrm{E}}U^{\phi^{j}})^{\phi^{k}}$. Then
$\lambda^{1+p^{i}}=\lambda^{p^{k}(1+p^{i})}.$
Note that $\lambda$ is a generator of ${\mathrm{E}}^{\times}$. Hence, we have that
@@
Hence, the characters of $g$ on $U\otimes_{\mathrm{E}}U^{\phi^{i}}$ and $(U\otimes_{\mathrm{E}}U^{\phi^{i}})^{\phi^{k}}$ cannot be equal. Therefore, $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ is a direct sum of non-isomorphic irreducible ${\mathbb{F}}_{p}G$-modules. ∎
Now we are ready to get the 6-th row of Table 1.
@@
Let $N=A_{p}(n,\theta)$ be defined in Example 2.4 such that $3\mid n$, $|\theta|=3$ and $p$ is an odd prime. Then $N$ is a $3$-orbit group and ${\mathrm{Aut}}(N)\cong p^{n^{2}}{:}{\Gamma\mathrm{L}}(3,p^{n/3})$.